# Entanglement transfer during quantum frequency conversion in gas-filled hollow-core fibers

### v1

@@ Abstract
Figure 3: Evolution of the photon numbers and concurrences along the fiber filled with 70 bar H2 and pumped with 115 $\mu$J pulse energy. (a) Photon numbers of the idler, $N_{I}$, mixing, $N_{M}$, and up-converted, $N_{U}$, frequency modes. The inset shows the simulated transverse intensity distribution (normalized to its maximum) of the fundamental core mode in a single-ring-type ARF similar to that used in Ref. Tyumenev2022 (see Table I at SM ). (b) Dynamics of the idler-mixing, $\mathcal{C}_{\text{I-M}}$, and idler-up-converted, $\mathcal{C}_{\text{I-U}}$, concurrences.
@@
In Fig. 3 (a), we show the photon numbers for the mixing, up-converted, and idler modes by blue solid, orange dashed, and green dashed-dotted lines, respectively. In Fig. 3 (b), we do the same for the idler-mixing and the idler-up-converted concurrences using red solid and purple dashed lines, respectively. We can observe that, until molecular coherence becomes significant at around the middle of the fiber, no substantial conversion dynamics occur; from that point on, the probability of frequency up-converting a mixing photon increases, and this is reflected in higher entanglement between idler and up-converted modes. Meanwhile, the idler-mixing concurrence decreases. Furthermore, notice that the crossings between the photon numbers and the concurrences occur at exactly the same point, around $z\approx$ 38 cm. This shows how closely related the transfer of entanglement and the transfer of photon population are. After a quarter of an oscillation, i.e. when $G_{U}|\xi|t=\pi/2$, we find the state $\left(|0,0,0\rangle+e^{-i\Omega t}|1,0,1\rangle\right)/\sqrt{2}$ for the idler, mixing, and up-converted, respectively. This state exemplifies that 100 % efficiency in the transfer of a photon from the mixing to the up-converted mode can be theoretically attainable, as can be seen from Eqs. (7) and (8). Note that our predictions could be tested in future experiments, since efforts to characterize the concurrence of mixed states have already been made Huang2009 , obtaining lower Mintert2007 and upper bounds Zhang2008 of this quantity.
@@
Figure 4: Concurrence as a function of the initial pump beam energy and the propagation distance $\bm{z}$. The plots show the evolution of the idler-mixing $\mathcal{C}_{\text{I-M}}$ and idler-up-converted $\mathcal{C}_{\text{I-U}}$ concurrences at a pressure of 70 bar. The white dashed lines correspond to the evolution displayed in Fig. 3.
In conclusion, we have explored the use of molecular modulation triggered by SRS in gas-filled ARFs for frequency up-conversion of entangled photon pairs, showing how entanglement can be efficiently transferred to the frequency-converted mode. To do so, we introduced a full quantum Hamiltonian description of the system capable of recovering the Maxwell-Bloch equations in the semiclassical limit. With it, we were able to characterize the state of the molecules as the vibrational coherence is established, and to use it in the analysis of the molecular modulation of quantum light states injected in H2-filled ARFs. We derived simple expressions governing the evolution of entanglement in the system that predict a close correlation with the evolution of the photon numbers. As the experiments in this area are rapidly progressing, we believe that this framework will be a useful resource for the design of novel fiber-based quantum transduction strategies that could be fully integrated with existing fiber networks, thereby bringing the dream of the future quantum networks one step closer.
@@ Acknowledgements.
We acknowledge financial support from HORIZON-CL4-2022-QUANTUM01-SGA project 101113946 OpenSuperQ-Plus100 of the EU Flagship on Quantum Technologies, the Spanish Ramón y Cajal Grant RYC-2020-030503-I, project Grants No. PID2021-125823NA-I00, PID2021-123131NA-I00, PID2021-122505OBC31 and TED2021-129959B-C21, funded by MICIU/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”, by “ERDF Invest in your Future”, by the “European Union NextGenerationEU/PRTR” and ”ESF+”. We also acknowledge support from the Spanish Ministry of Economic Affairs and Digital Transformation through the QUANTUM ENIA project call –Quantum Spain, and by the EU through the Recovery, Transformation and Resilience Plan– NextGenerationEU within the framework of the Digital Spain 2026 Agenda, as well as from the Basque Government through Grants No. IT1470-22 and IT1455-22 and ELKARTEK ($\mu$4Smart-KK-2023/00016, Ekohegaz II-KK-2023/00051, and KUBIT KK-2024/00105), and from the IKUR Strategy under the collaboration agreement between Ikerbasque Foundation and BCAM on behalf of the Department of Education of the Basque Government and the grant IKUR_IKA_23/04.
@@ I Experimental parameters
In this section, we present a list of the parameters extracted from the experiment in Ref. SM_Tyumenev2022 , which we have used to obtain the results presented in this manuscript.
@@
Table 1: Table containing experimental parameters from Ref. SM_Tyumenev2022 used in the simulations displayed in this letter.
@@
On the other hand, $\mathcal{V}$ is obtained by multiplying the total area of the fiber $A_{\text{fiber}}$ by the temporal width of the pulse $T_{\text{width}}$ and the speed of light $c$, i.e., $\mathcal{V}=cT_{\text{width}}A_{\text{fiber}}$. In this sense, $\mathcal{V}$ would follow the textbook definition of relevant volume that contains the energy of the radiation field, taking the volume integral of the electromagnetic energy density as reference. In order to provide a reasonable analytical expression for this volume, knowing that the real pulse has a finite duration, we define the temporal profile of the pulse as a piecewise function that is zero everywhere except in the relevant pulse region, where it would follow an amplitude distribution. In this regard, we consider the Bessel function $J_{0}$. Therefore, $T_{\text{width}}$ is estimated through the distance between the first zeros of $J_{0}$ that approximates the Gaussian distribution considered in Ref. SM_Tyumenev2022 for the pulse profile. The Bessel function used is determined by matching the integral of its square to the integral of the square of the previously mentioned Gaussian distribution in order to obtain the same total pulse energy while maintaining the same field amplitude factor. For the pulse transversal profile area, following the definition of volume containing the radiation, we considered $A_{\text{fiber}}$. This area is given by the inner diameter of the fiber at Ref. SM_Tyumenev2022 , excluding the capillaries. This is considered because, although hollow-core anti-resonant fibers present extremely low losses and provide tight modal confinement in the hollow region at the center of the multi-capillary microstructure, there is still some residual light intensity outside of the pulse’s effective mode area. With this approach, the value obtained for $\mathcal{V}$ leads to reasonable dynamics given the evolution time, given that $G_{S}$ and $G_{U}$ depend on the quantization volume.
In Fig. S1, we present the evolution of $\alpha_{P}$ and $\alpha_{S}$, the amplitudes of the coherent states that characterize the pump and the Stokes pulses inside the fiber, as studied in Ref. SM_Tyumenev2022 . In this study, molecular coherence inside a hollow-core fiber filled with hydrogen gas is developed through a stimulated Raman scattering process. A pump tone is used to excite the molecules, producing an increase on the population of the corresponding Stokes mode, as one can see in Fig. S1. In green, one can see the normalized amplitude of the pump field while, in red, we have the normalized amplitude of Stokes photons. Most of all dynamics occur at the latter half of the fiber, where coherence starts to develop.
@@ II From original to effective Hamiltonian
where $\omega_{i}$ is the frequency associated to the transition between vibrational levels $|i\rangle$ and $|i+1\rangle$ of the molecule, whereas $\omega_{l}$ represents the frequency of mode $l$ of the electric field, with $l\in\{P,S,A\}$ labelling the pump, Stokes, and anti-Stokes modes, respectively. On the other hand, we derive the interaction Hamiltonian, assuming that the interaction is dipolar, from the term $-\vec{\bf\mu}\vec{\bf E}(z)$. Here, $\vec{\bf\mu}$ and $\vec{\bf E}(z)$ are the dipole and the electric field operators, respectively, and can be written as
@@
$V_{I}(s)V_{I}(s^{\prime})$ $=$ $\sum_{i,j,k}\sum_{l,m}g_{i,j}^{l}g_{j,k}^{m}|i\rangle\langle k|\Big{(}a^{\dagger}_{l}a^{\dagger}_{m}e^{-i(\beta_{l}+\beta_{m})z}e^{i(\omega_{i}-\omega_{j}+\omega_{l})s}e^{i(\omega_{j}-\omega_{k}+\omega_{m})s}$
-- $a^{\dagger}_{l}a_{m}e^{-i(\beta_{l}-\beta_{m})z}e^{i(\omega_{i}-\omega_{j}+\omega_{l})s}e^{i(\omega_{j}-\omega_{k}-\omega_{m})s}-a_{l}a^{\dagger}_{m}e^{i(\beta_{l}-\beta_{m})z}e^{i(\omega_{i}-\omega_{j}-\omega_{l})s}e^{i(\omega_{j}-\omega_{k}+\omega_{m})s}$
$+$ $a_{l}a_{m}e^{i(\beta_{l}+\beta_{m})z}e^{i(\omega_{i}-\omega_{j}-\omega_{l})s}e^{i(\omega_{j}-\omega_{k}-\omega_{m})s}\Big{)}.$
@@
$\langle 1|V_{I}(s)\int_{0}^{s}\text{d}s^{\prime}\,V_{I}(s^{\prime})|1\rangle=i\sum_{k}|g_{1,k}^{l}|^{2}\sum_{l}a^{\dagger}_{l}a_{l}\left(\frac{1}{\Omega_{k,1}-\omega_{l}}+\frac{1}{\Omega_{k,1}+\omega_{l}}\right).$
@@
$H_{\text{eff}}$ $=$ $H_{0}+e^{-iH_{0}t/\hbar}H_{\text{eff}}^{I}(t)e^{iH_{0}t/\hbar}=\hbar\omega_{0}|0\rangle\langle 0|+\hbar\omega_{1}|1\rangle\langle 1|+\hbar\sum_{l}\omega_{l}a^{\dagger}_{l}a_{l}(|0\rangle\langle 0|+|1\rangle\langle 1|)+\hbar\sum_{l}a^{\dagger}_{l}a_{l}\left(\delta_{0,l}|0\rangle\langle 0|+\delta_{1,l}|1\rangle\langle 1|\right)$ (S32)
@@
$J^{\pm}\left|{\bf\frac{N}{2}},m_{z}\right\rangle$ $=$ $\sqrt{\frac{N}{2}\left(\frac{N}{2}+1\right)-m_{z}\left(m_{z}\pm 1\right)}\left|{\bf\frac{N}{2}},m_{z}\right\rangle.$ (S37)
@@
III Equations of motion
Going back to the single-molecule case, we want to show that we can recover the Maxwell-Bloch equations describing stimulated Raman scattering through the equations of motion associated to the effective Hamiltonian we derived in the previous section. In the following, we only describe the pump-Stokes interaction, considering that the anti-Stokes population is negligible throughout the process. We will work with the Hamiltonian
We first start in the Schrödinger picture, and go to an interaction picture with respect to the pump and Stokes fields. We transform our Hamiltonian by $\omega_{P}a^{\dagger}_{P}a_{P}+\omega_{S}a^{\dagger}_{S}a_{S}$ obtaining the Hamiltonian

### v2

@@ II.2 Mixing frequency conversion
Figure 3: Evolution of the average photon numbers and concurrences along the fiber filled with 70 bar H2 and pumped with 115 $\mu$J pulse energy. (a) Average photon numbers of the idler, $N_{I}$, mixing, $N_{M}$, and up-converted, $N_{U}$, frequency modes. The inset shows the simulated transverse intensity distribution (normalized to its maximum) of the fundamental core mode in a single-ring-type ARF similar to that used in Ref. Tyumenev2022, (see Table I in the supplementary material). Here, D represents the larger internal diameter of the fiber. (b) Dynamics of the idler-mixing, $C_{\text{I-M}}$, and idler-up-converted, $C_{\text{I-U}}$, concurrences.
Therefore, in this framework, the resulting equations for the evolution of the mixing, $N_{M}$, and the up-converted, $N_{U}$, average photon numbers during molecular modulation are
$N_{M}$ $=$ $\frac{1}{2}\cos^{2}\left(G_{U}|\xi|t\right),$ (10)
$N_{U}$ $=$ $\frac{1}{2}\sin^{2}\left(G_{U}|\xi|t\right).$ (11)
Meanwhile, we also study the dynamics of entanglement between the mixing and the up-converted frequencies through the concurrence Wootters1998 , an entanglement monotone used for bipartite mixed states. This is an appropriate choice in this case since we have entanglement between the idler and the mixing subsystems, but also between the idler and up-converted. Furthermore, the concurrence completely characterizes the entanglement of formation Hill1997 for a pair of two-level systems. In our case, we find that the idler-mixing and the idler-up-converted concurrences are
$C_{\text{I-M}}$ $=$ $\Big{|}\cos\left(G_{U}|\xi|t\right)\Big{|},$ (12)
$C_{\text{I-U}}$ $=$ $\Big{|}\sin\left(G_{U}|\xi|t\right)\Big{|}.$ (13)
Notice how entanglement transfer between mixing and up-converted modes is closely related to the evolution of the number of photons in each frequency mode, and thus higher conversion efficiencies lead to a more effective entanglement transfer. The time parameter $t$ here is considered to be related to the propagation distance $z$ inside the fiber used in Ref. Tyumenev2022, , $z\approx ct$. This is used to represent the evolution of the average photon numbers and the concurrences in Fig. 3, where the explicit time dependence of $\alpha_{P}$ and $\alpha_{S}$, as shown in the supplementary material, has been considered. These coefficients are obtained by numerically solving the semiclassical Maxwell-Bloch equations of motion for the pump and Stokes electric fieldsRaymer1990 ; Chen2024 . These already include the phenomenological $T_{2}$ factor accounting for the temporal dephasing of the molecular coherence. In addition, in order to obtain the coupling strength, we have estimated the quantization volume based on the geometric properties of the waveguide such as the larger internal diameter D (see Fig. 3) and the temporal length of the interaction.
III Results
@@
In Fig. 3 (a), we show the average photon numbers for the mixing, up-converted, and idler frequencies by blue solid, orange dashed, and green dashed-dotted lines, respectively. In Fig. 3 (b), we do the same for the idler-mixing and the idler-up-converted concurrences using red solid and purple dashed lines, respectively. We can observe that, until molecular coherence becomes significant at around the middle of the fiber, no substantial conversion dynamics occur; from that point on, the probability of frequency up-converting a mixing photon increases, and this is reflected in higher entanglement between idler and up-converted frequencies. Meanwhile, the idler-mixing concurrence decreases. Furthermore, notice that the crossings between the average photon numbers and the concurrences occur at exactly the same point, around $z\approx$ 38 cm, as a consequence of the close relation between the transfer of entanglement and the transfer of photon population. Note that, here, the maximum value for the average photon numbers is $1/2$ due to the type of photonic quantum state considered as input to the system. After a quarter of an oscillation, i.e. when $G_{U}|\xi|t=\pi/2$, we find the state $\left(|0,0,0\rangle+e^{-i\Omega t}|1,0,1\rangle\right)/\sqrt{2}$ for the idler, mixing, and up-converted, respectively. This state showcases that high efficiencies in the transfer of a photon from the mixing to the up-converted mode can be theoretically attainable, as can be seen from Eqs. (10) and (11). Note that our predictions could be tested in future experiments, since efforts to characterize the concurrence of mixed states have already been made Huang2009 , obtaining lower Mintert2007 and upper bounds Zhang2008 of this quantity.
@@
Figure 4: Concurrence as a function of the initial pump beam energy and the propagation distance $\bm{z}$. The plots show the evolution of the idler-mixing $C_{\text{I-M}}$ and idler-up-converted $C_{\text{I-U}}$ concurrences at a pressure of 70 bar. The white dashed lines correspond to the evolution displayed in Fig. 3.
IV Conclusion
In conclusion, we have explored the use of molecular modulation triggered by SRS in gas-filled ARFs for frequency up-conversion of entangled photon pairs, showing how entanglement can be efficiently transferred to the frequency-converted counterpart. To do so, we employ a quantum Hamiltonian description capable of recovering the Maxwell-Bloch equations in the semiclassical limit, as shown in the supplementary material. With it, we were able to characterize the state of the molecules as the vibrational coherence is established, and to use it in the analysis of the molecular modulation of quantum light states injected in gas-filled ARFs. We derived simple expressions governing the evolution of entanglement in the system, demonstrating the close relationship with the evolution of the average photon numbers.
We have found the single-mode treatment of both frequency and spatial modes to be satisfactory for the scope of this work, following the conditions of Ref. Tyumenev2022, . Nevertheless, a leap into a full multimode theory could be taken, for example, by following the quantum treatment in Ref. Wang2023, . This provides a description of the general effects of frequency translation in the shape of the second order time correlations of the experiment in Ref. Tyumenev2022, by considering the bandwidth in the quantum frequency mode distribution of the initial photonic state. Moreover, one can follow the works in Refs. Carman1970, ; Akhmanov1974, ; Raymer1979, ; Raymer1981, , which studied the role of the field bandwidths during the preparation of the Raman molecular coherence in the semiclassical regime. In addition, semiclassical treatments that deal with multiple spatial modes have also been considered Bauerschmidt2015_2 ; Agrawal2007 following the lines of Ref. Raymer1990, .
As the experiments in this area are rapidly progressing, we believe that this framework will be a useful resource for the design of novel fiber-based quantum transduction strategies that could be fully integrated with existing fiber networks, thereby bringing the dream of the future quantum networks one step closer.
Supplementary Material
In the supplementary material we provide more detailed calculations on the derivation of $H_{\text{eff}}$, the molecular state and the concurrences. The experimental parameters used at Ref. Tyumenev2022, and the subsequent time dependence of $\alpha_{P}$ and $\alpha_{S}$ can also be found.
@@ Acknowledgements.
We acknowledge financial support from HORIZON-CL4-2022-QUANTUM01-SGA project 101113946 OpenSuperQ-Plus100 of the EU Flagship on Quantum Technologies, the Spanish Ramón y Cajal Grant RYC-2020-030503-I, project Grants No. PID2021-125823NA-I00, PID2021-123131NA-I00, PID2021-122505OBC31 and TED2021-129959B-C21, funded by MICIU/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”, by “ERDF Invest in your Future”, by the “European Union NextGenerationEU/PRTR” and "ESF+", from the Basque Government through Grants No. IT1470-22 and IT1455-22 and ELKARTEK ($\mu$4Smart-KK-2023/00016, Ekohegaz II-KK-2023/00051, and KUBIT KK-2024/00105), and from the IKUR Strategy under the collaboration agreement between Ikerbasque Foundation and BCAM on behalf of the Department of Education of the Basque Government and the grant IKUR_IKA_23/04. ML acknowledges support from the predoctoral grant "Formación de Profesorado Universitario" FPU23/02350 from the Spanish Ministry of Science, Innovation and Universities (MICIU). This work has also been financially supported by the Ministry for Digital Transformation and the Civil Service of the Spanish Government through the QUANTUM ENIA project call – Quantum Spain project, and by the European Union through the Recovery, Transformation and Resilience Plan – NextGenerationEU within the framework of the Digital Spain 2026 Agenda.
Data Availability Statement
The data that support the findings of this study are available within the article and its supplementary material and from the corresponding author upon reasonable request.
Author Declarations
Conflict of Interest
The authors have no conflicts to disclose.
Author Contributions
T.G.-R. and A.M. contributed equally to this work.
Tasio Gonzalez-Raya: Formal analysis (equal); Methodology (lead); Software (equal); Writing – original draft (equal). Arturo Mena: Formal analysis (equal); Methodology (equal); Software (equal); Visualization (lead); Writing – original draft (equal). Miriam Lazo: Methodology (supporting). Luca Leggio: Visualization (supporting); Writing – review & editing (equal). David Novoa: Conceptualization (equal); Funding Acquisition (equal); Supervision (equal); Writing – original draft (supporting); Writing – review & editing (equal). Mikel Sanz: Conceptualization (lead); Methodology (equal); Funding Acquisition (equal); Supervision (equal); Writing – review & editing (equal);
@@ I Experimental parameters
In this section, we present a list of the parameters extracted from the experiment in Ref. SM_Tyumenev2022, (see Table 1), which we have used to obtain the results presented in this manuscript.
@@
Table 1: Table containing experimental parameters from Ref. SM_Tyumenev2022, used in the simulations displayed in this letter.
@@
On the other hand, $\mathcal{V}$ is obtained by multiplying the total area of the fiber $A_{\text{fiber}}$ by the temporal width of the pulse $T_{\text{width}}$ and the speed of light $c$, i.e., $\mathcal{V}=cT_{\text{width}}A_{\text{fiber}}$. In this sense, $\mathcal{V}$ would follow the textbook definition of relevant volume that contains the energy of the radiation field, taking the volume integral of the electromagnetic energy density as reference. In order to provide a reasonable analytical expression for this volume, knowing that the real pulse has a finite duration, we define the temporal profile of the pulse as a piecewise function that is zero everywhere except in the relevant pulse region, where it would follow an amplitude distribution. In this regard, we consider the Bessel function $J_{0}$. Therefore, $T_{\text{width}}$ is estimated through the distance between the first zeros of $J_{0}$ that approximates the Gaussian distribution considered in Ref. SM_Tyumenev2022, for the pulse profile. The Bessel function used is determined by matching the integral of its square to the integral of the square of the previously mentioned Gaussian distribution in order to obtain the same total pulse energy while maintaining the same field amplitude factor. For the pulse transversal profile area, following the definition of volume containing the radiation, we considered $A_{\text{fiber}}$. This area is given by the inner diameter of the fiber at Ref. SM_Tyumenev2022, , excluding the capillaries. This is considered because, although hollow-core anti-resonant fibers present extremely low losses and provide tight modal confinement in the hollow region at the center of the multi-capillary microstructure, there is still some residual light intensity outside of the pulse’s effective mode area. With this approach, the value obtained for $\mathcal{V}$ leads to reasonable dynamics given the evolution time, given that $G_{S}$ and $G_{U}$ depend on the quantization volume.
@@
In Fig. S1, we present the evolution of $\alpha_{P}$ and $\alpha_{S}$, the amplitudes of the coherent states that characterize the pump and the Stokes pulses inside the fiber, as studied in Ref. SM_Tyumenev2022, . In this study, molecular coherence inside a hollow-core fiber filled with hydrogen gas is developed through a stimulated Raman scattering process. A pump tone is used to excite the molecules, producing an increase on the population of the corresponding Stokes frequency mode, as one can see in Fig. S1. In green, one can see the normalized amplitude of the pump field while, in red, we have the normalized amplitude of Stokes photons. Most of all dynamics occur at the latter half of the fiber, where molecular coherence starts to develop.
@@ II From original to effective Hamiltonian
where $\omega_{i}$ is the frequency associated to the transition between vibrational levels $|i\rangle$ and $|i+1\rangle$ of the molecule, whereas $\omega_{l}$ represents the frequency of mode $l$ of the electric field, with $l\in\{P,S,A\}$ labelling the pump, Stokes, and anti-Stokes frequencies, respectively. On the other hand, we derive the interaction Hamiltonian, assuming that the interaction is dipolar, from the term $-\vec{\bf\mu}\vec{\bf E}(z)$. Here, $\vec{\bf\mu}$ and $\vec{\bf E}(z)$ are the dipole and the electric field operators, respectively, and can be written as
@@
$V_{I}(s)V_{I}(s^{\prime})$ $=$ $\sum_{i,j,k}\sum_{l,m}g_{i,j}^{l}g_{j,k}^{m}|i\rangle\langle k|\Big{(}a^{\dagger}_{l}a^{\dagger}_{m}e^{-i(\beta_{l}+\beta_{m})z}e^{i(\omega_{i}-\omega_{j}+\omega_{l})s}e^{i(\omega_{j}-\omega_{k}+\omega_{m})s^{\prime}}$
-- $a^{\dagger}_{l}a_{m}e^{-i(\beta_{l}-\beta_{m})z}e^{i(\omega_{i}-\omega_{j}+\omega_{l})s}e^{i(\omega_{j}-\omega_{k}-\omega_{m})s^{\prime}}-a_{l}a^{\dagger}_{m}e^{i(\beta_{l}-\beta_{m})z}e^{i(\omega_{i}-\omega_{j}-\omega_{l})s}e^{i(\omega_{j}-\omega_{k}+\omega_{m})s^{\prime}}$
$+$ $a_{l}a_{m}e^{i(\beta_{l}+\beta_{m})z}e^{i(\omega_{i}-\omega_{j}-\omega_{l})s}e^{i(\omega_{j}-\omega_{k}-\omega_{m})s^{\prime}}\Big{)}.$
@@
$\langle 1|V_{I}(s)\int_{0}^{s}\text{d}s^{\prime}\,V_{I}(s^{\prime})|1\rangle=i\sum_{k}|g_{1,k}^{l}|^{2}\sum_{l}\left(\frac{a^{\dagger}_{l}a_{l}}{\Omega_{k,1}-\omega_{l}}+\frac{a_{l}a^{\dagger}_{l}}{\Omega_{k,1}+\omega_{l}}\right).$
@@
$H_{\text{eff}}$ $=$ $H_{0}+e^{-iH_{0}t/\hbar}H_{\text{eff}}^{I}(t)e^{iH_{0}t/\hbar}=$ (S32)
$=$ $\hbar\omega_{0}|0\rangle\langle 0|+\hbar\omega_{1}|1\rangle\langle 1|+\hbar\sum_{l}\omega_{l}a^{\dagger}_{l}a_{l}(|0\rangle\langle 0|+|1\rangle\langle 1|)+\hbar\sum_{l}a^{\dagger}_{l}a_{l}\left(\delta_{0,l}|0\rangle\langle 0|+\delta_{1,l}|1\rangle\langle 1|\right)$
@@
$J^{\pm}\left|{\bf\frac{N}{2}},m_{z}\right\rangle$ $=$ $\sqrt{\frac{N}{2}\left(\frac{N}{2}+1\right)-m_{z}\left(m_{z}\pm 1\right)}\left|{\bf\frac{N}{2}},m_{z}\pm 1\right\rangle.$ (S37)
@@
As a final remark, we would like to point out that even though the adiabatic approximation performed through this second-order perturbation method is validated by semiclassical models SM_Raymer1990 , which accurately describes the results of the experiment in Ref. SM_Tyumenev2022, , issues can arise when the assumptions of ideal resonance or short evolution times are not met, which can lead to a non-Hermitian effective Hamiltonian. In Ref. SM_Begzjav2022, , different methods such as the Magnus expansion or the canonical transformation method are discussed, which can help surpass these issues of non-Hermiticity.
III Semiclassical approximation for photon modes
Using the Hamiltonian in Eq. (II) as a starting point, in the following we will consider that the anti-Stokes population is negligible, since the majority of the molecules remain in their ground state, and neglect the Stark shift terms. With these considerations, we will work with the following Hamiltonian, describing the pump-Stokes interactions.
$H_{\text{eff}}$ $=$ $\hbar\Omega J_{z}+\hbar\omega_{P}a^{\dagger}_{P}a_{P}+\hbar\omega_{S}a^{\dagger}_{S}a_{S}+\hbar\left(G_{S}e^{i\Delta\beta z}a_{P}a^{\dagger}_{S}J^{+}+G_{S}^{*}e^{-i\Delta\beta z}a^{\dagger}_{P}a_{S}J^{-}\right).$ (S39)
In the resonant case, $\Omega=\omega_{P}-\omega_{S}$, the interaction terms of the Hamiltonian, $\hbar\left(G_{S}e^{i\Delta\beta z}a_{P}a^{\dagger}_{S}J^{+}+G_{S}^{*}e^{-i\Delta\beta z}a^{\dagger}_{P}a_{S}J^{-}\right)$, commutes with the self energy terms, $\hbar\Omega J_{z}+\hbar\omega_{P}a^{\dagger}_{P}a_{P}+\hbar\omega_{S}a^{\dagger}_{S}a_{S}$, so we can study the evolution of the initial state under the Hamiltonian in the interaction picture with respect to these self energy terms, given in the equation below.
$H_{\text{eff}}^{I}=\hbar\left(G_{S}e^{i\Delta\beta z}a_{P}a^{\dagger}_{S}J^{+}+G_{S}^{*}e^{-i\Delta\beta z}a^{\dagger}_{P}a_{S}J^{-}\right).$ (S40)
In order to derive the molecular dynamics from this Hamiltonian, we will perform a semiclassical approximation in both photon modes. In the context of the experiment in Ref. SM_Tyumenev2022, , 115 $\mu$J pump pulses were used, which means that the initial pump state can be considered a coherent state with $\alpha_{P}\approx 2.48\times 10^{7}$. Then, since the annihilation of pump photons leads to the generation of intense laser radiation at the down-shifted Stokes frequency, we may also consider the Stokes field to be in a coherent state, with corresponding $\alpha_{S}$. With these considerations in mind, we define the following semiclassical Hamiltonian
$H_{\alpha}^{I}=\langle\alpha_{P}|\otimes\langle\alpha_{S}|\otimes\mathbb{1}H_{\text{eff}}^{I}|\alpha_{P}\rangle\otimes|\alpha_{S}\rangle\otimes\mathbb{1}=\hbar\left(G_{S}e^{i\Delta\beta z}\alpha_{P}\alpha_{S}^{*}J^{+}+G_{S}^{*}e^{-i\Delta\beta z}\alpha_{P}^{*}\alpha_{S}J^{-}\right),$ (S41)
which allows us to study the dynamics of the molecules. The Hamiltonian in Eq. (S41) is just the Hamiltonian obtained by replacing $a_{P}\rightarrow\alpha_{P}$, $a_{P}^{\dagger}\rightarrow\alpha_{P}^{*}$, $a_{S}\rightarrow\alpha_{S}$, $a_{S}^{\dagger}\rightarrow\alpha_{S}^{*}$ in in Eq. (S40). In order to illustrate the validity of this approximation, it would be useful to compare the evolution of the state of the molecules given by the semiclassical Hamiltonian in Eq. (S41) $H_{\alpha}^{I}$, to that obtained with the original Hamiltonian neglecting the Stark shift and the self-energy terms, $H_{\text{eff}}^{I}$, acting on the full state of both the pump and the Stokes modes and the molecules after taking the partial trace over the bosonic modes. The density matrix of the state of the molecules under the evolution of the Hamiltonian in Eq. (S41) (Eq. (4) in the manuscript) is given by
$\rho_{\alpha}=e^{-\frac{it}{\hbar}H_{\alpha}^{I}}|\Psi\rangle\langle\Psi|e^{\frac{it}{\hbar}H_{\alpha}^{I}}=|\Psi\rangle\langle\Psi|-\frac{it}{\hbar}\left[H_{\alpha}^{I},|\Psi\rangle\langle\Psi|\right]+\frac{1}{2}\left(\frac{it}{\hbar}\right)^{2}\left[H_{\alpha}^{I},\left[H_{\alpha}^{I},|\Psi\rangle\langle\Psi|\right]\right]+\ldots$ (S42)
where $|\Psi\rangle$ indicates the initial state of the molecules and we have expanded the exponential explicitly for the terms up to second order. On the other hand, the density matrix for the whole system is given by
$\rho=e^{-\frac{it}{\hbar}H_{\text{eff}}^{I}}|\alpha_{P}\rangle\otimes|\alpha_{S}\rangle\otimes|\Psi\rangle\langle\alpha_{P}|\otimes\langle\alpha_{S}|\otimes\langle\Psi|e^{\frac{it}{\hbar}H_{\text{eff}}^{I}}.$ (S43)
It is straightforward to see that if we expand the exponential and take partial trace over the photons states for the zeroth order term we would simply obtain $|\Psi\rangle\langle\Psi|$. We now examine the first order terms, $\rho^{(1)}$, that takes the following expression
$\rho^{(1)}=-\frac{it}{\hbar}\left[H_{\text{eff}}^{I},|\alpha_{P}\rangle\otimes|\alpha_{S}\rangle\otimes|\Psi\rangle\langle\alpha_{P}|\otimes\langle\alpha_{S}|\otimes\langle\Psi|\right].$ (S44)